1. calculate the wavelength in nm of the spectral line produced when an electron in a hydrogen atom undergoes the transition from the energy level n=6 to the level n=2

2. with what max kinetic energy will electrons be ejected when this metal is exposed to light with a wavelength of 245nm? threshold frequency is 2.65*10^14 s^-1
Answer #1

1)

1/lambda = 1.09737 x 10^7 x [1/2^2 - 1/6^2] = 2438600

Wavelength, lambda = 1/2438600 = 4.10 x 10^-7 m =

Wavelength, lambda = 410.07 nm

Comment in case any doubt please rate my answer ....

According to the HomeworkLib guidelines I can answer one question at a time so please post other questions seperately ...